### v1

@@ Abstract
Revolutionizing drug discovery demands more than just understanding molecular interactions—it requires generative models that can design novel ligands tailored to specific biological targets. While chemical Language Models (cLMs) have made strides in learning molecular properties, most fail to incorporate target-specific insights, restricting their ability to drive de-novo ligand generation. Chem42, a cutting-edge family of generative chemical Language Models, is designed to bridge this gap. By integrating atomic-level interactions with multimodal inputs from Prot42, a complementary protein Language Model [SNT+], Chem42 achieves a sophisticated cross-modal representation of molecular structures, interactions, and binding patterns. This innovative framework enables the creation of structurally valid, synthetically accessible ligands with enhanced target specificity. Evaluations across diverse protein targets confirm that Chem42 surpasses existing approaches in chemical validity, target-aware design, and predicted binding affinity. By reducing the search space of viable drug candidates, Chem42 could accelerate the drug discovery pipeline, offering a powerful generative AI tool for precision medicine. The Chem42 models, available at huggingface.co/inceptionai, set a new benchmark in molecule property prediction, conditional molecule generation, and target-aware ligand design.
Figure 1: Example of a Chem42-generated ligand designed to bind the Y220C mutant p53 protein (PDB ID: 8a32), shown on the right (panel (c)). For comparison, panel (a) presents the binding pose of the recently proposed small-molecule stabilizer JC769 as estimated using DiffDock-L [CDF+24], while panel (b) displays the stabilizer JC769 as originally reported in the literature [SCDD+22]. The protein target context is captured using our protein Language Model (Prot42 [SNT+]) and integrated into the chemical Language Model (Chem42) to guide ligand generation. Calculated metrics—including Drug-likeness, Synthetic Accessibility, and Binding Affinity—demonstrate the effectiveness of Chem42 in achieving state-of-the-art molecular design results.
@@ 1 Introduction
Drug discovery is inherently multimodal, since ligand activity is dictated not only by the structure of a molecule, but also by its complex interactions with protein targets. To address this limitation, we introduce Chem42, a generative chemical-protein modeling framework that integrates our protein Language Model (pLM), named Prot42 [SNT+], with our chemical Language Model for target-aware ligand generation. This cross-modal approach enhances molecular design by embedding the protein context into the generative process, enabling the model to generate valid, biologically relevant, and target-specific ligands. Figure 1 illustrates how Chem42 generates a ligand for the Y220C p53 mutant protein (a tumor suppressor called also ”Guardian of the Genome”), comparing its binding pose to JC769, a known stabilizer (PDB ID: 8a32) [SCDD+22]. Both ligands bind to the same pocket, demonstrating the model’s capacity to design structure-guided molecules. Given a target protein such as p53, we prompt Prot42 with its sequence, embedding its molecular features into Chem42 via cross-attention, allowing for ligand generation guided by target-specific interactions. To the best of our knowledge, this is the first framework where a cLM and a pLM collaborate to explicitly account for protein context in ligand generation, marking a significant advancement in AI-driven drug design.
@@ 3 Methodology
We introduce a novel approach leveraging our family of chemical Language Models (cLMs), specifically designed for drug-like and target-aware ligand generation. Unlike conventional molecular representations, our approach explicitly captures atomic-level interactions, providing a chemically meaningful and information-rich representation of molecules. By embedding molecules within a language that inherently models atomic interactions, we establish a precise and generative framework for designing compounds with desired pharmacological properties. Our contributions in this paper are threefold: (i) We develop a novel chemical language that encodes atomic interactions, enhancing both the representation and generation of drug-like molecules; (ii) We introduce a multi-omics modality integration framework by combining our chemical language model with a protein Language Model (Prot42 [SNT+]), enabling target-aware ligand generation aligned with structural and functional properties of the target; and (iii) We conduct a comprehensive evaluation of our approach in terms of chemical validity, binding affinity prediction, and target specificity. By bridging chemical and protein Language Models, we advance toward a rational and data-driven drug design paradigm, unlocking new possibilities for highly specific therapeutic development. Our chemical language models (cLMs) were pre-trained on the extensive UniChem dataset [CDG+13] using a decoder-only, autoregressive LLaMA-inspired architecture [TMS+23]. To determine the optimal model size and hyperparameters, we employed maximal update parametrization ($\mu$P) [mup-mutransfer], leveraging a 38M-parameter proxy model. After fine-tuning hyperparameters through this proxy, we conducted full-scale pre-training, ensuring both scalability and efficiency. Evaluations demonstrated that a target tokens-per-parameter (TPP) ratio of 50 provided the optimal balance, achieving the lowest validation loss while maintaining computational efficiency. This approach optimizes performance while reducing computational overhead, setting a new benchmark for efficient and scalable chemical modeling.
@@ 3.4 Protein-binding Ligand Generation
Given a dataset $\mathcal{D}$ of protein-ligand pairs, our objective is to generate ligand sequences $\hat{S}_{l}$ that exhibit high affinity toward a target protein. Let $(S_{p},S_{l})\in\mathcal{D}$ represent a protein-ligand pair, where $S_{p}=(s_{1},s_{2},...,s_{m})$ is a sequence of $m$ amino acid tokens, and $S_{l}=(l_{1},l_{2},...,l_{n})$ is a sequence of $n$ ligand tokens, each belonging to a predefined vocabulary $\mathcal{V}_{l}$. The protein model (Prot42 [SNT+] $M_{p}$ encodes $S_{p}$ into a sequence of hidden representations $\mathbf{E}_{p}=(e_{1},e_{2},...,e_{m})$, where each $e_{i}\in\mathbb{R}^{2048}$ captures structural and functional properties of the protein. Similarly, the ligand model (Chem42) $M_{l}$ encodes $S_{l}$ into a sequence of latent embeddings $\mathbf{H}_{l}=(h_{1},h_{2},...,h_{n})$, where each $h_{t}\in\mathbb{R}^{1280}$ represents the hidden state of the chemical model. To integrate protein context into ligand generation, protein embeddings are projected into the hidden dimension of the ligand model using a learnable transformation $\mathbf{E}_{p}^{\prime}=\theta_{p}\mathbf{E}_{p}$, where $\theta_{p}\in\mathbb{R}^{2048\times 1280}$. These transformed protein embeddings are incorporated into ligand representations using a cross-attention mechanism. Attention scores are computed as $A=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{1280}}\right)$, where the query, key, and value projections are defined as $Q=\theta_{q}\mathbf{H}_{l},\quad K=\theta_{k}\mathbf{E}_{p}^{\prime},\quad V=\theta_{v}\mathbf{E}_{p}^{\prime},$ with learnable weight matrices $\theta_{q},\theta_{k},\theta_{v}\in\mathbb{R}^{1280\times 1280}$. The resulting protein-conditioned ligand representations are given by $\mathbf{C}_{l}=AV$, which are combined with the original ligand embeddings as $\mathbf{H}_{l}^{\prime}=\mathbf{H}_{l}+\mathbf{C}_{l}$. These transformed representations $\mathbf{H}_{l}^{\prime}$ are then passed through the CFM decoder layers. The generation of ligands follows an autoregressive process in which each token $\hat{l}_{t}$ is predicted based on previously generated tokens and protein-informed features. The probability distribution over the ligand vocabulary $\mathcal{V}_{l}$ is given by $y_{t}=\text{Softmax}(\theta_{h}h_{t}^{\prime}+\theta_{c}C_{t}),$ where $\theta_{h},\theta_{c}$ are learnable parameters, and the next ligand token is selected as $\hat{l}_{t}=\arg\max_{j}(y_{t}^{j}).$ This process continues until a termination token is reached. The generated ligand sequence $\hat{S}_{l}=(\hat{l}_{1},\hat{l}_{2},...,\hat{l}_{\hat{n}})$ is then ranked using a scoring function $R(\hat{S}_{l})$, which evaluates molecular validity, drug-likeness (QED) and synthetic accessibility (SA). Thus, the entire protein-ligand generation process can be represented as: $\hat{S}_{l}=M_{l}(\mathbf{E}_{p}^{\prime},\mathbf{H}_{l}),$ where the parameters of this model are $\theta=\{\theta_{p},\theta_{q},\theta_{k},\theta_{v},\theta_{h},\theta_{c}\}$.
@@ 4.5 Target-aware Ligand Generation
This task focuses on generating drug-like molecules that can bind to target proteins with unknown binding sites. We used the PDBBind 2020 training set [^4] https://huggingface.co/datasets/jglaser/binding_affinity, which contains more than 16,000 unique protein-ligand pairs. For docking simulations, we employ DiffDock [CDF+24], and for computing the binding affinity, we use Prodigy-Lig [VSK+19]. Our approach is evaluated on nine target proteins;
@@
These targets were not part of the model’s training data. For each target protein, we generate 10 SMILES sequences, evaluating their drug-likeness (QED), synthetic accessibility (SA), both computed using RDKit [BHF+20]. From these, we select the best molecule based on the properties mentioned, perform docking simulations with DiffDock, and finally calculate the binding score. Table 9 presents a comparative analysis of our multi-omics modality protein-ligand generation model against PMN [NH24] across multiple target proteins. Our models consistently outperform PMN in QED and SA, demonstrating improved molecular quality and ease of synthesis. Note that we compare absolute binding affinity values, as their magnitude directly reflects interaction strength, ensuring a consistent assessment of binding stability across different targets. Figure 5 visualizes some of our generated ligands.
Target ID Drug-likeness QED ($\uparrow$) Synthetic Accessibility SA ($\downarrow$) BA ($\downarrow$) kcal/mol
PMN [NH24] Base Chem42 PMN Base Chem42 PMN Base Chem42
8a32 - 0.8738 0.9217 - 2.3787 1.8490 - 6.92 7.42
1iep 0.322 0.8523 0.8617 7.609 2.5813 2.5944 9.946 6.8 7.69
2rgp 0.428 0.7328 0.8965 3.391 2.4189 3.1475 11.936 7.94 9.68
3eml 0.584 0.8909 0.8408 7.268 2.2339 3.0985 25.939 7.93 8.38
3ny8 0.807 0.7006 0.9439 5.99 2.4527 2.5785 11.257 10.35 8.37
4rlu 0.479 0.8650 0.7696 2.979 2.3596 3.6555 11.250 7.87 6.63
4unn 0.161 0.8701 0.9026 4.567 2.5952 1.8674 10.752 8.33 8.57
5mo4 0.432 0.9044 0.8156 6.330 3.1990 3.1051 11.812 8.73 7.67
7l11 0.136 0.8200 0.8911 7.912 3.4333 3.6877 11.220 9.26 8.09
2nzt - 0.8771 0.9024 - 2.5347 3.2397 - 7.01 7.45
Table 9: Comparison of Baseline and Chem42 metrics (QED, SA, BA) for different PDB IDs - first two metrics are calculated by RDkit, BA are calculated using AutoDock Vina (as Vina score in kcal mol-1) for PMN [NH24] and using Prodigy-Lig [VSK+19] for ours. Bold: best performance; Green indicates best performance within the Chem42 and Baseline models.
@@ 5 Conclusion
In this work, we introduced Chem42, a new family of chemical Language Models (CLMs) designed for target-aware ligand generation by integrating atomic interaction modeling with protein Language Models (Prot42). Our approach addresses a critical limitation of existing molecular generative models by incorporating target-specific information through a multimodal framework, enabling the generation of ligands that are both chemically valid and biologically relevant. Through extensive evaluation, we demonstrated that Chem42 outperforms existing CLMs in terms of molecular validity, target specificity, and predicted binding affinity. The ability of our model to capture atomic-level interactions enhances the precision of ligand design, while integration with protein embeddings ensures context-aware molecular generation, even in cases where binding sites are not well defined. Our experimental results validate the effectiveness of this approach, showing improved drug-likeness (QED), synthetic accessibility (SA), and binding affinity (BA) compared to baseline methods.
The broader impact of this work extends beyond ligand generation. By bridging chemical and protein representations, we have pave the way for more accurate and efficient AI-driven drug discovery pipelines. Future directions include further scaling of the model, optimizing protein-ligand docking accuracy, and extending multimodal learning to incorporate structural and biochemical constraints. By integrating chemistry-aware and biology-aware representations, our work advances toward a data-driven, rational approach to drug design, ultimately accelerating the discovery of novel therapeutics with high specificity and efficacy.
@@ Acknowledgment
The authors are thankful to the DiffDock team for the release of the DiffDock-L [CDF+24], a state-of-the-art model for molecular docking. All ligand-protein docking results reported in this paper are obtained using DiffDock-L. All protein visualizations reported in this paper are obtained using the UCSF ChimeraX software.

### v2

@@ Abstract
Revolutionizing drug discovery demands more than just understanding molecular interactions—it requires generative models that can design novel ligands tailored to specific biological targets. While chemical Language Models (cLMs) have made strides in learning molecular properties, most fail to incorporate target-specific insights, restricting their ability to drive de-novo ligand generation. Chem42, a cutting-edge family of generative chemical Language Models, is designed to bridge this gap. By integrating atomic-level interactions with multimodal inputs from Prot42, a complementary protein Language Model [STN+25], Chem42 achieves a sophisticated cross-modal representation of molecular structures, interactions, and binding patterns. This innovative framework enables the creation of structurally valid, synthetically accessible ligands with enhanced target specificity. Evaluations across diverse protein targets confirm that Chem42 surpasses existing approaches in chemical validity, target-aware design, and predicted binding affinity. By reducing the search space of viable drug candidates, Chem42 could accelerate the drug discovery pipeline, offering a powerful generative AI tool for precision medicine. The Chem42 models, available at huggingface.co/inceptionai, set a new benchmark in molecule property prediction, conditional molecule generation, and target-aware ligand design.
Figure 1: Example of a Chem42-generated ligand designed to bind the Y220C mutant p53 protein (PDB ID: 8a32), shown on the right (panel (c)). For comparison, panel (a) presents the binding pose of the recently proposed small-molecule stabilizer JC769 as estimated using DiffDock-L [CDF+24], while panel (b) displays the stabilizer JC769 as originally reported in the literature [SCDD+22]. The protein target context is captured using our protein Language Model (Prot42 [STN+25]) and integrated into the chemical Language Model (Chem42) to guide ligand generation. Calculated metrics—including Drug-likeness, Synthetic Accessibility, and Binding Affinity—demonstrate the effectiveness of Chem42 in achieving state-of-the-art molecular design results.
@@ 1 Introduction
Drug discovery is inherently multimodal, since ligand activity is dictated not only by the structure of a molecule, but also by its complex interactions with protein targets. To address this limitation, we introduce Chem42, a generative chemical-protein modeling framework that integrates our protein Language Model (pLM), named Prot42 [STN+25], with our chemical Language Model for target-aware ligand generation. This cross-modal approach enhances molecular design by embedding the protein context into the generative process, enabling the model to generate valid, biologically relevant, and target-specific ligands. Figure 1 illustrates how Chem42 generates a ligand for the Y220C p53 mutant protein (a tumor suppressor called also ”Guardian of the Genome”), comparing its binding pose to JC769, a known stabilizer (PDB ID: 8a32) [SCDD+22]. Both ligands bind to the same pocket, demonstrating the model’s capacity to design structure-guided molecules. Given a target protein such as p53, we prompt Prot42 with its sequence, embedding its molecular features into Chem42 via cross-attention, allowing for ligand generation guided by target-specific interactions. To the best of our knowledge, this is the first framework where a cLM and a pLM collaborate to explicitly account for protein context in ligand generation, marking a significant advancement in AI-driven drug design.
@@ 3 Methodology
We introduce a novel approach leveraging our family of chemical Language Models (cLMs), specifically designed for drug-like and target-aware ligand generation. Unlike conventional molecular representations, our approach explicitly captures atomic-level interactions, providing a chemically meaningful and information-rich representation of molecules. By embedding molecules within a language that inherently models atomic interactions, we establish a precise and generative framework for designing compounds with desired pharmacological properties. Our contributions in this paper are threefold: (i) We develop a novel chemical language that encodes atomic interactions, enhancing both the representation and generation of drug-like molecules; (ii) We introduce a multi-omics modality integration framework by combining our chemical language model with a protein Language Model (Prot42 [STN+25]), enabling target-aware ligand generation aligned with structural and functional properties of the target; and (iii) We conduct a comprehensive evaluation of our approach in terms of chemical validity, binding affinity prediction, and target specificity. By bridging chemical and protein Language Models, we advance toward a rational and data-driven drug design paradigm, unlocking new possibilities for highly specific therapeutic development. Our chemical language models (cLMs) were pre-trained on the extensive UniChem dataset [CDG+13] using a decoder-only, autoregressive LLaMA-inspired architecture [TMS+23]. To determine the optimal model size and hyperparameters, we employed maximal update parametrization ($\mu$P) [mup-mutransfer], leveraging a 38M-parameter proxy model. After fine-tuning hyperparameters through this proxy, we conducted full-scale pre-training, ensuring both scalability and efficiency. Evaluations demonstrated that a target tokens-per-parameter (TPP) ratio of 50 provided the optimal balance, achieving the lowest validation loss while maintaining computational efficiency. This approach optimizes performance while reducing computational overhead, setting a new benchmark for efficient and scalable chemical modeling.
@@ 3.4 Protein-binding Ligand Generation
Given a dataset $\mathcal{D}$ of protein-ligand pairs, our objective is to generate ligand sequences $\hat{S}_{l}$ that exhibit high affinity toward a target protein. Let $(S_{p},S_{l})\in\mathcal{D}$ represent a protein-ligand pair, where $S_{p}=(s_{1},s_{2},...,s_{m})$ is a sequence of $m$ amino acid tokens, and $S_{l}=(l_{1},l_{2},...,l_{n})$ is a sequence of $n$ ligand tokens, each belonging to a predefined vocabulary $\mathcal{V}_{l}$. The protein model (Prot42 [STN+25] $M_{p}$ encodes $S_{p}$ into a sequence of hidden representations $\mathbf{E}_{p}=(e_{1},e_{2},...,e_{m})$, where each $e_{i}\in\mathbb{R}^{2048}$ captures structural and functional properties of the protein. Similarly, the ligand model (Chem42) $M_{l}$ encodes $S_{l}$ into a sequence of latent embeddings $\mathbf{H}_{l}=(h_{1},h_{2},...,h_{n})$, where each $h_{t}\in\mathbb{R}^{1280}$ represents the hidden state of the chemical model. To integrate protein context into ligand generation, protein embeddings are projected into the hidden dimension of the ligand model using a learnable transformation $\mathbf{E}_{p}^{\prime}=\theta_{p}\mathbf{E}_{p}$, where $\theta_{p}\in\mathbb{R}^{2048\times 1280}$. These transformed protein embeddings are incorporated into ligand representations using a cross-attention mechanism. Attention scores are computed as $A=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{1280}}\right)$, where the query, key, and value projections are defined as $Q=\theta_{q}\mathbf{H}_{l},\quad K=\theta_{k}\mathbf{E}_{p}^{\prime},\quad V=\theta_{v}\mathbf{E}_{p}^{\prime},$ with learnable weight matrices $\theta_{q},\theta_{k},\theta_{v}\in\mathbb{R}^{1280\times 1280}$. The resulting protein-conditioned ligand representations are given by $\mathbf{C}_{l}=AV$, which are combined with the original ligand embeddings as $\mathbf{H}_{l}^{\prime}=\mathbf{H}_{l}+\mathbf{C}_{l}$. These transformed representations $\mathbf{H}_{l}^{\prime}$ are then passed through the CFM decoder layers. The generation of ligands follows an autoregressive process in which each token $\hat{l}_{t}$ is predicted based on previously generated tokens and protein-informed features. The probability distribution over the ligand vocabulary $\mathcal{V}_{l}$ is given by $y_{t}=\text{Softmax}(\theta_{h}h_{t}^{\prime}+\theta_{c}C_{t}),$ where $\theta_{h},\theta_{c}$ are learnable parameters, and the next ligand token is selected as $\hat{l}_{t}=\arg\max_{j}(y_{t}^{j}).$ This process continues until a termination token is reached. The generated ligand sequence $\hat{S}_{l}=(\hat{l}_{1},\hat{l}_{2},...,\hat{l}_{\hat{n}})$ is then ranked using a scoring function $R(\hat{S}_{l})$, which evaluates molecular validity, drug-likeness (QED) and synthetic accessibility (SA). Thus, the entire protein-ligand generation process can be represented as: $\hat{S}_{l}=M_{l}(\mathbf{E}_{p}^{\prime},\mathbf{H}_{l}),$ where the parameters of this model are $\theta=\{\theta_{p},\theta_{q},\theta_{k},\theta_{v},\theta_{h},\theta_{c}\}$.
@@ 4.5 Target-aware Ligand Generation
This task focuses on generating drug-like molecules that can bind to target proteins with unknown binding sites. We used the PDBBind 2020 training set [^4] https://huggingface.co/datasets/jglaser/binding_affinity, which contains more than 16,000 unique protein-ligand pairs. For docking simulations and computing the binding affinity, we used AutoDock-Vina [TO10]. Our approach is evaluated on eight target proteins;
@@
These targets were not part of the model’s training data. For each target protein, we generate 500 SMILES sequences, evaluating their drug-likeness (QED), synthetic accessibility (SA), both computed using RDKit [BHF+20]. From these, we select the best molecule based on the properties mentioned, perform docking simulations with AutoDock Vina [TO10], and finally get the binding affinities. Table 9 presents a comparative analysis of our multi-omics modality protein-ligand generation model against PMN [NH24] across multiple target proteins. While binding affinities are generally comparable between the two models—with each showing strengths on different targets—our models consistently achieve higher QED and lower SA scores, demonstrating improved molecular quality and ease of synthesis-key factors for successful downstream drug development. Figure 5 visualizes some of our generated ligands.
Table 9: Performance comparison between PMN, Chem42 (507M), and Chem42 (597M) on eight protein targets. Higher QED and more negative BA are desirable; lower SA indicates easier synthesis. Bold indicates the best performance per metric per target.
Target QED ($\uparrow$) SA ($\downarrow$) BA ($\downarrow$) (kcal/mol) PMN Chem42 (507M) Chem42 (597M) PMN Chem42 (507M) Chem42 (597M) PMN Chem42 (507M) Chem42 (597M) 1iep 0.322 0.613 0.360 7.609 4.125 3.152 -9.946 -10.860 -10.35 2rgp 0.428 0.526 0.562 3.391 2.195 2.334 -11.936 -10.690 -10.10 3eml 0.584 0.624 0.534 7.268 1.895 3.271 -25.939 -10.590 -10.40 3ny8 0.807 0.383 0.397 5.990 2.583 3.250 -11.257 -12.070 -10.70 4rlu 0.479 0.854 0.803 2.979 1.735 1.765 -11.250 -9.100 -9.80 4unn 0.161 0.333 0.463 2.323 2.331 3.628 -10.752 -10.190 -10.40 5mo4 0.432 0.782 0.415 6.330 2.321 2.753 -11.812 -10.260 -10.23 7l11 0.136 0.513 0.645 7.912 2.978 2.540 -11.220 -7.580 -8.43
@@ 5 Conclusion
In this work, we introduced Chem42, a new family of chemical Language Models (CLMs) designed for target-aware ligand generation by integrating atomic interaction modeling with protein Language Models (Prot42). Our approach addresses a critical limitation of existing molecular generative models by incorporating target-specific information through a multimodal framework, enabling the generation of ligands that are both chemically valid and biologically relevant. Through extensive evaluation, we demonstrated that Chem42 outperforms existing CLMs in terms of molecular validity, target specificity, and is comparable in predicting binding affinity. The ability of our model to capture atomic-level interactions enhances the precision of ligand design, while integration with protein embeddings ensures context-aware molecular generation, even in cases where binding sites are not well defined. Our experimental results validate the effectiveness of this approach, showing improved drug-likeness (QED), synthetic accessibility (SA), and on par with binding affinity (BA) in comparison to baseline methods.
The broader impact of this work extends beyond ligand generation. By bridging chemical and protein representations, we have paved the way for more accurate and efficient AI-driven drug discovery pipelines. Future directions include further scaling of the model, optimizing protein-ligand docking accuracy, and extending multimodal learning to incorporate structural and biochemical constraints. By integrating chemistry-aware and biology-aware representations, our work advances toward a data-driven, rational approach to drug design, ultimately accelerating the discovery of novel therapeutics with high specificity and efficacy.
@@ Acknowledgment
The authors are thankful to the AutoDock Vina team [TO10], a state-of-the-art model for molecular docking. All ligand-protein docking results reported in this paper are obtained using this tool. All protein visualizations reported in this paper were obtained using the UCSF ChimeraX software.